Hankins Corporation has 8.1 million shares of common stock outstanding, 300,000 shares of 4.1 percent preferred stock outstanding, par value of $100; and 185,000 bonds with a semiannual coupon rate of 5.5 percent outstanding, par value $2,000 each. The common stock currently sells for $57 per share and has a beta of 1.15, the preferred stock has a par value of $100 and currently sells for $99 per share, and the bonds have 18 years to maturity and sell for 107 percent of par. The market risk premium is 6.6 percent, T-bills are yielding 3.3 percent, and the company’s tax rate is 24 percent.

A. What is the firm’s market value capital structure?

B. If the company is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows?

Solve for:

A. Debt

Preferred Stock

Equity

B. Discount Rate
Answer #1

A. Market Value of Shares =Number of shares*Share Price =8100000*57 =461,700,000
Market Value of Preferred Shares =Number of Shares*Preferred Share Price =300000*99 =29,700,000
Market Value of Debt =Price of Bonds*Number of Bonds =107%*2000*185000 =395,900,000
Total Market Value =461,700,000+29,700,000+395,900,000=887,300,000
Weight of Equity =461,700,000/887,300,000 =52.0343% or 52.03%
Weight of Preferred Stock =29,700,000/887,300,000 =3.3472% or 3.35%
Weight of Debt =29,700,000/887,300,000 =44.6185% or 44.62%

b) Cost of equity =Risk Free rate+Beta*(Market return-Risk Free rate) =3.3%+1.15*6.6% =10.89%
Cost of Preferred stock =Dividend /price =4.1%*100/99 =4.14%
Coupon =2000*5.5%/2 =55
Number of Periods =18*2 =36
Price of Bond =107%*2000 =2140
YTM using excel formula =2*RATE(36,55,-2140,2000)=4.49082%
Cost of Capital or Discount Rate =Weight of Equity*Cost of Equity+weight of Preferred Stock*Cost of Preferred Stock+Weight of Debt*Cost of Debt*(1-Tax Rate) =52.0343%*10.89%+3.3472%*4.14%+44.6185%*4.49082%*(1-24%) =7.47%